OpenStudy (anonymous):

What is the Square root of -72?

OpenStudy (anonymous):

@ganeshie8 @girlwithsquirrels14 @Ashleyisakitty

ganeshie8 (ganeshie8):

72 = 36x2

ganeshie8 (ganeshie8):

\(\large \sqrt{-72} \) \(\large \sqrt{-36 \times 2} \)

OpenStudy (anonymous):

-6 sqrt 2 6 sqrt -2 6sqrt 2i 6i sqrt 2

ganeshie8 (ganeshie8):

\(\large \sqrt{-72} \) \(\large \sqrt{-36 \times 2} \) \(\large \sqrt{-1 \times 36 \times 2} \)

OpenStudy (anonymous):

Those are the options

OpenStudy (anonymous):

-72 = 8.4852814

ganeshie8 (ganeshie8):

u want ansurr or learn how to do this ? :)

OpenStudy (anonymous):

Learn how please

ganeshie8 (ganeshie8):

u sure ? dont lie

OpenStudy (anonymous):

Yeah I want to understand!

ganeshie8 (ganeshie8):

pull the 36 out, it becomes 6 ?

ganeshie8 (ganeshie8):

good :) \(\large \sqrt{-72} \) \(\large \sqrt{-36 \times 2} \) \(\large \sqrt{-1 \times 36 \times 2} \) \(\large 6\sqrt{-1 \times 2} \)

ganeshie8 (ganeshie8):

see if above steps make sense

OpenStudy (anonymous):

Eh not really ! I dont know this

ganeshie8 (ganeshie8):

\(\large \sqrt{-72} \) \(\large \sqrt{-36 \times 2} \)

ganeshie8 (ganeshie8):

till that step u okays ? :)

OpenStudy (anonymous):

yess!

ganeshie8 (ganeshie8):

since, sqrt(36) = 6 it becomes :- \(\large 6 \sqrt{-1 \times 2} \)

ganeshie8 (ganeshie8):

remember this for the rest of ur course time : sqrt(-1) = i so it becomes :- \(\large 6 i \sqrt{2} \)

OpenStudy (anonymous):

Oh okay . And what are the steps for solving 2 divided by 2 + 5i?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

multiply top and bottom wid conjugate

ganeshie8 (ganeshie8):

\(\large \frac{2}{2+5i} \) multiply top and bottom wid conjugate 2-5i \(\large \frac{2}{2+5i} \times \frac{2-5i}{2-5i}\)

OpenStudy (anonymous):

wait so 4 and 10i

ganeshie8 (ganeshie8):

\(\large \frac{2}{2+5i} \) multiply top and bottom wid conjugate 2-5i \(\large \frac{2}{2+5i} \times \frac{2-5i}{2-5i}\) \(\large \frac{2(2-5i)}{(2+5i)(2-5i)}\) since, \((a+ib)(a-ib) = a^2+b2, \) \(\large \frac{2(2-5i)}{2^2+5^2}\) \(\large \frac{4-10i}{29}\)

OpenStudy (anonymous):

Wow thats tricky for me ..I see what you mean tho . thank you !

ganeshie8 (ganeshie8):

glad u cud make some sense of that :) yw !